# Long and Short-Term Constraints Driven Safe Reinforcement Learning for Autonomous Driving

### v1

@@ Abstract
Reinforcement learning (RL) has been widely used in decision-making tasks, but it cannot guarantee the agent’s safety in the training process due to the requirements of interaction with the environment, which seriously limits its industrial applications such as autonomous driving. Safe RL methods are developed to handle this issue by constraining the expected safety violation costs as a training objective, but they still permit unsafe state occurrence, which is unacceptable in autonomous driving tasks. Moreover, these methods are difficult to achieve a balance between the cost and return expectations, which leads to learning performance degradation for the algorithms. In this paper, we propose a novel algorithm based on the long and short-term constraints (LSTC) for safe RL. The short-term constraint aims to guarantee the short-term state safety that the vehicle explores, while the long-term constraint ensures the overall safety of the vehicle throughout the decision-making process. In addition, we develop a safe RL method with dual-constraint optimization based on the Lagrange multiplier to optimize the training process for end-to-end autonomous driving. Comprehensive experiments were conducted on the MetaDrive simulator. Experimental results demonstrate that the proposed method achieves higher safety in continuous state and action tasks, and exhibits higher exploration performance in long-distance decision-making tasks compared with state-of-the-art methods.
@@ I Introduction
Autonomous driving (AD) is rapidly developing, and great achievements have been made [1, 2, 3]. As a new technology, end-to-end autonomous driving attracted researchers’ attention in recent years [4, 5, 6]. In comparison to traditional pipeline-based methods [7], it simplifies the architecture of AD system and replaces the multi-module system with a single end-to-end model, where raw sensor data are directly input to generate corresponding driving commands such as the steering angle, braking, and throttle.
In recent years, reinforcement learning (RL) has become the mainstream framework for end-to-end autonomous driving [8, 9], since it does not require a large number of samples like imitation learning. The vehicle agents can be trained by interacting with the environment without human intervention and adjust their behaviors based on the feedback provided by the environment. However, this trial-and-error property brings potential safety risks for training an effective model in a practical end-to-end AD system.
@@
Safe reinforcement learning, which is often modeled as a constrained Markov decision process (CMDP) [10], is developed to handle the safety issue of traditional RL methods by introducing a constraint in the training process. In these methods, a cost is usually predefined to evaluate the risk of the agent at current time step. The generalized constraint emphasizes that the expected cost should be less than a predetermined threshold. Many safe RL methods [11], [12], [13], [14], [15], [16] use the expected cost, similar to the expected reward, as a secondary objective which mainly includes the descriptions of decision risks or losses, which is a long-term objective constraint. These methods exhibit good safety performance, convergence, and stability of network training in many intelligent control tasks like obstacle avoidance. However, they only guarantee the safety after the training convergence, but do not consider the state safety of the agent in training process, which may lead to serious accidents when applying in autonomous driving. Moreover, using the expected cost as a constraint is a modification of the goal-oriented learning approach that is similar to reward shaping, which may lead to inconsistency in the optimal strategy and result in performance reduction.
In this paper, we propose a novel safe RL method with long and short-term constraints (LSTC) for decision-making tasks in end-to-end autonomous driving, as shown in Fig.1. The long-term constraint is designed based on the cost signals, with the goal of keeping the vehicle safe throughout a complete task. The short-term constraint, where a learnable forward-domain safety model checker that is designed to verify the state safety of a trajectory, aims to guarantee the short-term state safety that the vehicle explores. To solve the optimization problem for the proposed LSTC, we also develop a safe RL method with dual-constraint optimization based on the Lagrange multiplier.
In conclusion, the main contributions are as follows.
  • We propose the long and short-term constraints for safe RL in this paper, which aims to address the issue of unrestricted exploration in traditional CMDP-based methods and enhance the safety of vehicle agents.
  • To ensure the safety without sacrificing performance, we develop a safe RL method with dual-constraint optimization for autonomous driving, which can solve the constrained optimization problem of the proposed LSTC in the training process.
@@ II-A End-to-end Autonomous Driving
Traditional pipeline-based autonomous driving has achieved remarkable success in recent years [17, 18], but it remains some limitations including the difficulty of high computational costs and cumulative errors between modules. End-to-end methods can handle these issues by replacing the complex modules [19, 20] with a single deep neural network model [21, 22, 23], and they are often categorized into two types: imitation learning (IL) and reinforcement learning (RL) [24].
@@
IL-based methods train an agent to learn the optimal strategy by imitating the expert demonstration. One of the most important and widely used methods is the behavior cloning (BC) [25]. Codevilla et al. [26] proposed the Conditional Imitation Learning (CIL) model, which inputs monocular images, self-vehicle speed measurements, high-level driving decision-making and directly predicts the vehicle’s lateral and longitudinal controls. Many researchers expand on [26] by incorporating additional input like global routes, positional data, or point clouds [27],[28],[29]. Due to the richer input data, these methods exhibit impressive adaptability and resilience across a wide array of scenarios. BC demonstrates proficiency when dealing with states within the training distribution. However, the generalization ability to new states diminishes due to the cumulative errors in actions, which limits its industrial applications. In addition, IL-based methods require a large number of labeled data for training, on which many approaches [30], [26], [31] rely to achieve good performance. However, it’s challenging to make the collected data cover all the possible states in complex urban driving scenarios.
@@
RL-based methods gradually attract researchers’ attention in recent years due to the advantage of not requiring a large number of labeled data. The agent in RL learns to make decisions by trial and error in the process of interacting with the environment, in which it receives the feedback in the form of rewards or penalties for its actions and aims to maximize the total cumulative reward. Alizadeh et al. [32] train a agent with the deep Q network (DQN) to make high-level decisions. In real-world scenarios, Kendall et al. [33] use the deep deterministic policy gradient (DDPG) algorithm, and the vehicle can drive in simple environments like human drivers. [34], [35], [36], [37] indicate that the Proximal Policy Optimization (PPO) algorithm performs well in the tasks with continuous driving actions. However, these methods are developed in simulated environments since they ignore the safety problem in the training and testing processes.
@@
Safety consideration is increasingly being emphasized, especially for AD tasks in the real environment. Introducing safety checker is a common approach, which can recognize whether an action belongs to the set of safe actions. Zhang et al. [30] propose SafeDAgger, where the safety policy can predict whether the decision made by the primary policy is safe. If the decision is predicted to be unsafe, the reference policy is activated to take over. The reference policy serves as the final safeguard to determine the safety of the ultimate decision. Bouton et al. [38] design a probability model to predict the probability of each decision and compare this probability with a threshold to identify safe actions, enabling the vehicle to safely reach the destination. The probability model exhibits high accuracy in discrete state-action spaces, but it is not applicable in continuous high-dimensional spaces. Inspired by the idea of safety checker, we focus on the long short-term safety of decisions and propose a safe RL method for continuous state-action spaces in AD, which can validate the safety of forward state trajectories of the vehicle agent.
@@ II-B Safe reinforcement learning
The aforementioned methods provide solutions for constrained optimization problems and have been proven to be effective. However, they overlook the safety of the agent during the training process, essentially requiring trial and error to minimize errors. In contrast, our approach ensures the safety of the agent during both training and deployment through a dual-constraint mechanism.
III Preliminaries
@@
A feasible state $s_{f}$ indicates that there is no risk associated with this state, while an infeasible state $s_{inf}$ suggests that a vehicle in this state may encounter collisions with other vehicles, obstacles, buildings, etc. They are defined based on predefined rules. We define the set composed of all $s_{f}$ as the feasible state space $\mathcal{S}_{f}$ and the set composed of all $s_{inf}$ as the infeasible state space $\mathcal{S}_{inf}$, as shown in Fig. 1.
@@
The state trajectory $\tau^{n}$ with total $n$ time steps, which is generated from state $s$ following policy $\pi$, is defined as Eq.1.
$\begin{split}\tau^{n}&:=(s_{t}|s_{t}=s,\pi)\\ &=\left\{s_{t},s_{t+1},s_{t+2},s_{t+3},...,s_{t+n}|s_{t}=s,\pi\right\},\end{split}$ (1)
@@ IV Methodology
CMDP-based methods use the cost value function as the constraint, which lacks state-based safety and is potentially dangerous for AD vehicles. In this paper, we aim for the vehicle to drive within a reliable region during both exploration and deployment. For this purpose, we extend the state-based constraint by using forward state trajectories for the safety verification and propose our safe RL method with LSTC in order to constrain the vehicle’s exploration in the safe space and ensure the vehicle safety during the training process.
@@ IV-A Long and short-term constraints
where $c_{t+1}$ is the cost value from the environment at the time step $t+1$, which is calculated by the predefined cost function like the reward function in RL methods. The details of calculating the reward and cost at a time step will be shown in Sec. V-A. The cost value function $V_{\pi}^{c}(s)$ is used to estimate the long-term cost that the agent can expect when executing a certain policy in different states, so we refer to this constraint as the long-term constraint. Since the proposed long-term constraint is designed to modify the objective function so as to decrease the final cost value, ensuring the agent safety from the perspective of the long-term learning objective.
Short-term constraint: AD vehicles need to consistently drive within the feasible state space $\mathcal{S}_{f}$ for safety. Although the proposed long-term constraint can reduce the risk after training convergence, it struggles to handle the tasks of requiring the state safety such as autonomous driving. Nevertheless, as defined in Def. 3, our objective policy is to select the optimal policy within the feasible space, where a feasible policy can ensure that the agent is safe in every state under that policy since ensuring the safety in each state is particularly crucial for AD vehicles.
In order to handle this issue, we decompose the driving task from the initial state to the terminal state into several short finite-length state trajectories and aim to ensure the agent safety in the decomposed trajectories, for which we need to make all the trajectories to fall within the feasible state space, as shown in Fig. 1. In other words, each of $n$-step state trajectory $\tau^{n}$ starting from state $s$ needs to be within the feasible space $\mathcal{S}_{f}$ for all $s$ based on a feasible policy $\pi$. Accordance to this idea, we develop the short-term constraint based on the state, as formulated by Eq. 7.
$\begin{split}\tau^{n}&:=(s_{t}|s_{0}=s,\pi)\in\mathcal{S}_{f}.\end{split}$ (7)
@@
$\begin{split}&\theta\leftarrow\theta+\alpha_{\theta}\nabla_{\theta}L(\theta,\lambda_{s},\lambda_{l}),\\ &\lambda_{s}\leftarrow\lambda_{s}-\alpha_{s}\nabla_{\lambda_{s}}L(\theta,\lambda_{s},\lambda_{l}),\\ &\lambda_{l}\leftarrow\lambda_{l}-\alpha_{l}\nabla_{\lambda_{l}}L(\theta,\lambda_{s},\lambda_{l}),\end{split}$ (11)
@@ IV-B Safe RL method with dual-constraint optimization
The value network and cost value network optimize the parameters to reduce the error in the value function, thereby improving the estimation of cumulative reward and cost. We use the mean squared error (MSE) as their loss functions, as shown by Eqs. 15 and 16, respectively.
$\begin{split}L_{C_{r}}(\omega_{r})=\mathbbm{E}[(\hat{V}_{\omega_{r}}(s)-V_{\omega_{r}}(s))^{2}],\end{split}$ (15)
$\begin{split}L_{C_{c}}(\omega_{c})=\mathbbm{E}[(\hat{V}_{\omega_{c}}(s)-V_{\omega_{c}}(s))^{2}],\end{split}$ (16)
where $L_{C_{r}}(\omega_{r})$ and $L_{C_{c}}(\omega_{c})$ are the loss functions of the value network and the cost value network, respectively. $\hat{V}_{\omega_{r}}(s)$ and $\hat{V}_{\omega_{c}}(s)$ are the target value function and the target cost value function, respectively.
@@
                   Sample from environment and store data in $\mathcal{D}$
@@
       Sample batch from $\mathcal{D}$, First update the Lagrange Multiplier,
@@
      Update the validation network: $\phi\leftarrow\phi+\alpha_{\phi}\nabla_{\phi}L(\phi)$
       Update the policy network: $\theta\leftarrow\theta+\alpha_{\theta}\nabla_{\theta}L(\theta,\lambda,\phi)$
       Update the value network: $\omega_{r}\leftarrow\omega_{r}+\alpha_{\omega_{r}}\nabla_{\omega}L_{\omega_{r}}(\omega_{r})$
       Update the cost-value network: $\omega_{c}\leftarrow\omega_{c}+\alpha_{\omega_{c}}\nabla_{\omega}L_{\omega_{c}}(\omega_{c})$

### v2

@@ Abstract
Reinforcement learning (RL) has been widely used in decision-making and control tasks, but the risk is very high for the agent in the training process due to the requirements of interaction with the environment, which seriously limits its industrial applications such as autonomous driving systems. Safe RL methods are developed to handle this issue by constraining the expected safety violation costs as a training objective, but the occurring probability of an unsafe state is still high, which is unacceptable in autonomous driving tasks. Moreover, these methods are difficult to achieve a balance between the cost and return expectations, which leads to learning performance degradation for the algorithms. In this paper, we propose a novel algorithm based on the long and short-term constraints (LSTC) for safe RL. The short-term constraint aims to enhance the short-term state safety that the vehicle explores, while the long-term constraint enhances the overall safety of the vehicle throughout the decision-making process, both of which are jointly used to enhance the vehicle safety in the training process. In addition, we develop a safe RL method with dual-constraint optimization based on the Lagrange multiplier to optimize the training process for end-to-end autonomous driving. Comprehensive experiments were conducted on the MetaDrive simulator. Experimental results demonstrate that the proposed method achieves higher safety in continuous state and action tasks, and exhibits higher exploration performance in long-distance decision-making tasks compared with state-of-the-art methods.
@@ I Introduction
Autonomous driving (AD) is rapidly developing, and great achievements have been made in recent years [1, 2, 3, 4, 5]. As a new technology, end-to-end autonomous driving, which learns driving policies from collected data or environmental interaction rather than pre-defined complex rules, attracted researchers’ attention [6, 7, 8]. In comparison to traditional pipeline-based methods [9, 10], it simplifies the architecture of AD systems and replaces the multi-module system with a single end-to-end model, where raw sensor data are directly input to generate corresponding driving commands such as the steering angle, braking, and throttle.
In recent years, reinforcement learning (RL) has become the mainstream framework for end-to-end autonomous driving [11] since it does not require a large number of labeled samples like imitation learning. The vehicle agents can be trained by interacting with the environment without human intervention and adjust their behaviors based on the feedback provided by the environment. However, this trial-and-error property brings potential safety risks for training an effective model in a practical end-to-end AD system.
@@
Safe reinforcement learning, which is often modeled as a constrained Markov decision process (CMDP) [12], is developed to handle the safety issue of traditional RL methods by introducing a constraint in the training process. In these methods, a cost is usually predefined to evaluate the risk of the agent at current time step. The generalized constraint emphasizes that the expected cost should be less than a predetermined threshold. Many safe RL methods [13, 14, 15, 16, 17, 18] use the expected cost, similar to the expected reward, as a secondary objective which mainly includes the descriptions of decision risks or losses, which is a long-term objective constraint. These methods exhibit good safety performance, convergence, and stability of network training in many intelligent control tasks. However, they only guarantee the safety after the training convergence, but do not consider the state safety of the agent in training process, which may lead to serious accidents when applying in autonomous driving. Moreover, using the expected cost as a constraint is a modification of the goal-oriented learning approach that is similar to reward shaping, which may lead to inconsistency in the optimal strategy and result in performance reduction.
In this paper, we propose a novel safe RL method with Long and Short-Term Constraints (LSTC) for decision-making tasks in end-to-end autonomous driving, as shown in Fig.1. The long-term constraint is designed based on the cost signals, with the goal of keeping the vehicle safe throughout a complete task. The short-term constraint, where a learnable forward-domain safety model checker that is designed to verify the state safety of a trajectory, aims to enhance the short-term state safety that the vehicle explores. To solve the optimization problem for the proposed LSTC, we also develop a safe RL method with dual-constraint optimization based on the Lagrange multiplier. In conclusion, the main contributions are as follows.
  • We propose the long and short-term constraints for safe RL, which aims to address the issue of unrestricted exploration in existing RL-based methods and enhance the agent safety in the training process.
  • To improve safety performance without sacrificing learning performance, we develop a safe RL method with dual-constraint optimization for autonomous driving, which can solve the constrained optimization problem of the proposed LSTC.
@@ II-A End-to-end Autonomous Driving
Traditional pipeline-based autonomous driving has achieved remarkable success in recent years [19, 20], but it remains some limitations including the difficulty of high computational costs and cumulative errors between modules [21]. End-to-end methods can handle these issues by replacing the complex modules with a single deep neural network model [22], and they are often categorized into two types: imitation learning (IL) and reinforcement learning (RL) [6].
@@
IL-based methods train an agent to learn the optimal strategy by imitating the expert demonstration. One of the most important and widely used methods is the behavior cloning (BC) [23]. Codevilla et al. [24] proposed the Conditional Imitation Learning (CIL) model, which inputs monocular images, self-vehicle speed measurements, high-level driving decision-making and directly predicts the vehicle’s lateral and longitudinal controls. Many researchers expand on [24] by incorporating additional input like global routes, positional data, or point clouds [25, 26, 27]. Due to the richer input data, these methods exhibit impressive adaptability and resilience across a wide array of scenarios. BC demonstrates proficiency when dealing with states within the training distribution. However, the generalization ability to new states diminishes due to the cumulative errors in actions, which limits its industrial applications. In addition, IL-based methods require a large number of labeled data for training, on which many approaches [28, 24, 29] rely to achieve good performance. However, it’s challenging to make the collected data cover all the possible states in complex urban driving scenarios.
@@
RL-based methods [30, 31] gradually attract researchers’ attention in recent years due to the advantage of not requiring a large number of labeled data. The agent in RL learns to make decisions by trial and error in the process of interacting with the environment, in which it receives the feedback in the form of rewards or penalties for its actions and aims to maximize the total cumulative reward. Alizadeh et al. [32] train a agent with the deep Q network (DQN) to make high-level decisions. In real-world scenarios, Kendall et al. [33] use the deep deterministic policy gradient (DDPG) algorithm, and the vehicle can drive in simple environments like human drivers. [34, 35, 36, 37] indicate that the Proximal Policy Optimization (PPO) algorithm performs well in the tasks with continuous driving actions. However, these methods are developed in simulated environments since they ignore the safety problem in the training and testing processes.
@@
Safety consideration is increasingly being emphasized, especially for AD tasks in the real environment. Introducing safety checker is a common approach, which can recognize whether an action belongs to the set of safe actions. Zhang et al. [28] propose SafeDAgger, where the safety policy can predict whether the decision made by the primary policy is safe. If the decision is predicted to be unsafe, the reference policy is activated to take over. The reference policy serves as the final safeguard to determine the safety of the ultimate decision. Bouton et al. [38] design a probability model to predict the probability of each decision and compare this probability with a threshold to identify safe actions, enabling the vehicle to safely reach the destination. The probability model exhibits high accuracy in discrete state-action spaces, but it is not applicable in continuous high-dimensional spaces. Inspired by the idea of safety checker, we focus on the long and short-term safety of decisions and propose a safe RL method for continuous state-action spaces in AD, which can validate the safety of forward state trajectories of the vehicle agent.
@@ II-B Safe reinforcement learning
The aforementioned methods provide solutions for constrained optimization problems and have been proven to be effective. However, they overlook the state safety of the agent during the training process, essentially requiring trial and error to minimize errors. In contrast, our approach enhances the agent safety during both training and deployment through a dual-constraint mechanism.
III Preliminaries and problem formulation
@@
In our method, the whole state space of the vehicle is divided into two parts: feasible state space $\mathcal{S}_{f}$ and infeasible state space $\mathcal{S}_{inf}$, as shown in Fig. 1. A feasible state $s_{f}\in\mathcal{S}_{f}$ indicates that there is no risk associated with this state, while the infeasible state $s_{inf}\in\mathcal{S}_{inf}$ suggests that a vehicle in this state may encounter collisions with other vehicles, obstacles, buildings, etc. We differentiate them through specific rules, including speed limits, time-to-collision (TTC), and heading angle deviation, etc. following [43].
@@
The state trajectory $\tau^{n}$ with total $n$ time steps, which is generated from state $s_{t}$ following policy $\pi$, is defined as Eq.1.
$\begin{split}\tau^{n}&=\left\{s_{t}|\pi\right\}_{t}^{t+n}=\left\{s_{t},s_{t+1},s_{t+2},s_{t+3},...,s_{t+n}|\pi\right\},\end{split}$ (1)
@@ IV Methodology
CMDP-based methods use the cost value function as the constraint, which lacks state-based safety and is potentially dangerous for AD vehicles. In this paper, we aim for the vehicle to drive within a reliable region during both exploration and deployment. For this purpose, we extend the state-based constraint by using forward state trajectories for the safety verification and propose our safe RL method with LSTC in order to constrain the vehicle’s exploration in the safe space and enhance the vehicle safety during the training process.
@@ IV-A Long and short-term constraints
where $c_{t+1}$ is the cost value from the environment at the time step $t+1$, which is calculated by the predefined cost function like the reward function in RL methods. The details of calculating the reward and cost at a time step will be shown in Sec. V-A. The cost value function $V_{\pi}^{c}(s)$ is used to estimate the long-term cost that the agent can expect when executing a certain policy in different states, so we refer to this constraint as the long-term constraint. Since the proposed long-term constraint is designed to modify the objective function so as to decrease the final cost value, it can enhance the agent safety from the perspective of the long-term learning objective.
Short-term constraint: AD vehicles need to consistently drive within the feasible state space $\mathcal{S}_{f}$ for safety. Although the proposed long-term constraint can reduce the risk after training convergence, it struggles to handle the tasks of requiring the state safety such as autonomous driving. Nevertheless, as defined in Def. 3, our objective policy is to select the optimal policy within the feasible space, where a feasible policy can ensure that AD vehicles are safe in every state under that policy since the safety in each state is particularly crucial for AD vehicles.
In order to address this issue, we decompose the driving process from the initialization to termination into sequential states, and expect each state along with several subsequent states in the vehicle’s trajectory to fall in the feasible state space, which aims to ensure that the AD vehicle always navigates in the feasible state space and meets safety requirements, as illustrated in Fig. 1. In other words, each $n$-step state trajectory $\tau^{n}$ starting from the state $s$ needs to adhere to a feasible policy $\pi$ to remain in the feasible space $\mathcal{S}_{f}$. According to this idea, we develop the short-term constraint based on each state safety of the AD vehicle, as formulated in Eq. 7.
$\begin{split}\tau^{n}=\left\{s_{t}|\pi\right\}_{t}^{t+n}\in\mathcal{S}_{f}.\end{split}$ (7)
@@
$\left\{\begin{array}[]{lr}\theta\leftarrow\theta+\alpha_{\theta}\nabla_{\theta}L(\theta,\lambda_{s},\lambda_{l}),&\\ \lambda_{s}\leftarrow\lambda_{s}-\alpha_{s}\nabla_{\lambda_{s}}L(\theta,\lambda_{s},\lambda_{l}),&\\ \lambda_{l}\leftarrow\lambda_{l}-\alpha_{l}\nabla_{\lambda_{l}}L(\theta,\lambda_{s},\lambda_{l}),&\end{array}\right.$ (11)
@@ IV-B Safe RL method with dual-constraint optimization
The value network and cost value network optimize the parameters through minimizing the temporal-difference (TD) error, thereby improving the estimation of cumulative reward and cost. We use the mean squared error (MSE) as their loss functions, as shown by Eqs. 15 and 16, respectively.
$\begin{split}L_{C_{r}}(\omega_{r})=\mathbbm{E}[(r_{t+1}+\gamma V_{\omega_{r}}(s_{t+1})-V_{\omega_{r}}(s))^{2}],\end{split}$ (15)
$\begin{split}L_{C_{c}}(\omega_{c})=\mathbbm{E}[(c_{t+1}+\gamma V_{\omega_{c}}(s_{t+1})-V_{\omega_{c}}(s))^{2}],\end{split}$ (16)
where $L_{C_{r}}(\omega_{r})$ and $L_{C_{c}}(\omega_{c})$ are the loss functions of the value network and the cost value network, respectively. $r_{t+1}+\gamma V_{\omega_{r}}(s_{t+1})$ and $c_{t+1}+\gamma V_{\omega_{c}}(s_{t+1})$ are the TD targets. $V_{\omega_{r}}(s)$ and $V_{\omega_{c}}(s)$ are the estimations of the value function and the cost value function, respectively.
@@
                   Sample from environment and store $(s,a,r)$ in $\mathcal{D}$
@@
       Sample minibatch from $\mathcal{D}$, First update the Lagrange Multiplier,
@@
      Update the validation network by Eq. 19: $\phi\leftarrow\phi+\alpha_{\phi}\nabla_{\phi}L(\phi)$
       Update the policy network by Eq. 12: $\theta\leftarrow\theta+\alpha_{\theta}\nabla_{\theta}L(\theta,\lambda,\phi)$
       Update the value network by Eq. 15: $\omega_{r}\leftarrow\omega_{r}+\alpha_{\omega_{r}}\nabla_{\omega}L_{\omega_{r}}(\omega_{r})$
       Update the cost-value network Eq. 16: $\omega_{c}\leftarrow\omega_{c}+\alpha_{\omega_{c}}\nabla_{\omega}L_{\omega_{c}}(\omega_{c})$